OpenStudy (anonymous):

fifth term of a geometric sequence is 112 and the ninth term is 1,792. Find the common ratio of the sequence

OpenStudy (campbell_st):

well you'll solve 2 equations by subsitution 6th term \[112=ar^{(6-1)}\] 9th term \[1792 = ar^{9-1}\] rewrite the equation for the 9th term in terms of the 6th term \[1792 = ar^5 \times r^3\] substitute the 6th term and then solve for r hope it helps